# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ B.3 Proof of part (c)
$d_{uv}^{*}(\rho)=\sqrt{\dfrac{n}{n+\tau^{-2}}}d_{uv}(\rho)R_{uv}+\sqrt{n}\left(\dfrac{\rho^{2}\sigma_{C,v}\sigma_{C,u}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}}\right).$
Due to pseudo-posterior concentration of $\sigma_{C,j}^{2}$ to $\sigma_{0j}^{2}$ and convergence in probability of $||\mu_{j}||_{2}^{2}$ to $||\lambda_{0j}||_{2}^{2}$, we have pseudo-posterior concentration of $d_{uv}(\rho)$ around 00 as $n\to\infty$ for any finite $\rho>0$. Furthermore, with probability at least $1-o_{P_{0}}(1)$,
$\sqrt{n}\left(\dfrac{{\sigma}_{C,u}{\sigma}_{C,v}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}}\right)\lesssim\dfrac{1}{\sqrt{n}}.$
Thus, the pseudo-posterior of $d_{uv}^{*}(\rho)$ concentrates around 00 asymptotically. As $n\to\infty$, Lemma F.2 leads to
[…]
For $u=v$, we first observe that $\sqrt{n}(\Psi_{C,uu}-T_{uu})=\sqrt{n}(L_{C,uu}-\|\mu_{u}\|_{2}^{2})+\sqrt{n}(\sigma_{C,u}^{2}-\delta_{u}^{2}).$ For a given $1\leq j\leq p$, using the decomposition in part (b) of the proof of Theorem 3.6, we can express
$\sqrt{n}(\sigma_{C,j}^{2}-\delta_{j}^{2})=-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2}+\mathcal{G}_{j},$
for all $1\leq j\leq p$, where $|\mathcal{G}_{j}|\lesssim(1/\sqrt{n})+(\sqrt{n}/p_{n})$ with probability at least $1-o_{P_{0}}(1).$ Letting $\mathcal{N}_{j}=-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2},$ we have $\mathcal{N}_{j}/\sqrt{2\sigma_{0u}^{4}}\implies N(0,1)$, using a normal approximation to the Gamma distribution. Thus, we can write
$\sqrt{n}\left(\Psi_{C,uu}-T_{uu}\right)=\left(2\rho\,\sigma_{0u}\|\lambda_{0u}\|_{2}\right)R_{uu}+\mathcal{N}_{u}+\mathcal{G}_{u}.$
@@
and $f_{n}=n/(n+\tau^{-2})^{2}.$ Since $y^{(u)}=M_{0}\lambda_{0u}+\epsilon^{(u)}$, we can decompose $S_{uv}$ as
@@
Observe that $f_{n}=(1/n)+\Delta_{n}$, where $|\Delta_{n}|\asymp 1/n^{2}$. We first break up $S_{uv}$ into two parts $S_{uv}=D_{uv}+R_{uv},$ where
@@
We first deal with $R_{uv}$ and show that $\sqrt{n}R_{uv}=o_{P_{0}}(1)$. We use throughout that $\|M_{0}\|\asymp\sqrt{n}$ and $\|\epsilon^{(u)}\|_{2}\lesssim\sqrt{n}$ with probability at least $1-o(1),$ using Lemma E.8. First consider bounding $\sqrt{n}f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}$ as follows:
@@
The first term is $\sqrt{n}f_{n}\left|\epsilon^{(u)}U_{0}U_{0}^{\top}\epsilon^{(v)}\right|\,=O_{P_{0}}(\sqrt{n}f_{n})=O_{P_{0}}(1/\sqrt{n})$, since $U_{0}^{\top}\epsilon^{(u)}\sim N(0,\sigma_{0u}^{2}\mathbb{I}_{k})$ and $k$ is finite. The second term is handled by observing that
@@
    We first observe that with probability at least $1-o(1)$,
@@
    $\leq\sqrt{n}\Delta_{n}\|M_{0}^{\top}M_{0}\|||\lambda_{0u}||_{2}||\lambda_{0v}||_{2}\lesssim\dfrac{1}{\sqrt{n}}.$
@@
    We next observe that with probability at least $1-o(1),$
@@
$\begin{split}\mathcal{\xi}_{0,uv}^{2}=\mbox{var}\left[(\lambda_{0u}^{\top}\eta_{01})(\lambda_{0v}^{\top}\eta_{01})\right]&=\left\{\begin{array}[]{ll}(\lambda_{0u}^{\top}\lambda_{0v})^{2}+||\lambda_{0u}||_{2}^{2}||\lambda_{0v}||_{2}^{2},&\text{for $u\neq v$},\\ 2\|\lambda_{0u}\|_{2}^{4},&\text{for $u=v$.}\end{array}\right.\end{split}$
@@
where $Z_{uv}\sim N(0,1)$ and $Z_{uv}$ is independent of $M_{0}.$ Let
$l_{0,uv}^{2}=\sigma_{0v}^{2}||\lambda_{0u}||_{2}^{2}+\sigma_{0u}^{2}||\lambda_{0v}||_{2}^{2}.$
@@
Since $\|M_{0}\|\asymp\sqrt{n}$, we have $\left|l_{uv}(M_{0})+l_{0,uv}\right|=O_{P_{0}}(1)$ and $\left|l_{uv}(M_{0})+l_{0,uv}\right|\geq l_{0,uv}\geq 2c_{1}c_{2}>0$, from Assumptions 3.2 and 3.3. Thus, the third term can be handled by observing
@@
$=O_{P_{0}}\left(\dfrac{1}{\sqrt{n}}\right),$
as $\left\|\dfrac{M_{0}^{\top}M_{0}}{n}-\mathbb{I}_{k}\right\|=O_{P_{0}}\left(\dfrac{1}{\sqrt{n}}\right)$ from Lemma E.1. Since $Z_{uv}$ is independent of $M_{0}$, Lemma F.1 immediately implies that
[…]
@@
since $\|AB\|\geq\|A\|s_{\min}(B)$ for any matrices $A,B$, with $s_{min}(B)$ denoting the smallest singular value of $B$. Both $\Lambda_{0}$ and $M_{0}$ have exactly $k$ positive singular values since the rank is $k$. Thus, we have
$\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}U_{0}U_{0}^{\top}X_{0}\|\geq s_{\min}^{2}(\Lambda_{0})s_{\min}^{2}(M_{0})\|UU^{\top}-U_{0}U_{0}^{\top}\|.$
Since the Schatten $\infty$-norm is identical to the spectral norm, Theorem 2 in Luo et al., (2021) implies
@@
Thus with probability at least $1-o(1)$,
@@
Using Lemma E.8, we have $s_{\min}(M_{0})\asymp\|M_{0}\|\ \asymp\sqrt{n}$ with probability at least $1-o(1)$, since $k=o(n)$. Under Assumption 3.2, $s_{\min}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$. Suppose $\sigma_{\text{sum}}^{2}=\sum_{j=1}^{p}\sigma_{0j}^{2}$ and $\sigma_{\max}^{2}=\max_{1\leq j\leq p}\sigma_{0j}^{2}=\mathcal{O}(1)$ by Assumption 3.3. Lemma E.2 implies $\|E\|\lesssim(\sigma_{\text{sum}}+\sqrt{n}\sigma_{\max})$ with probability at least $1-o(1)$. Since $\sigma_{\text{sum}}\leq\sqrt{p_{n}}\sigma_{\max}$, one has $\|E\|\lesssim\sqrt{n}+\sqrt{p_{n}}$ with probability at least $1-o(1)$, as $n\to\infty$. Thus, with probability at least $1-o(1)$,
@@ E.2 Other Lemmas for Theorem 1
$\mathbb{E}\|E_{0}^{\top}E_{0}-n_{1}\mathbb{I}_{n_{2}}\|\leq\left(n_{1}^{1/2}+n_{2}^{1/2}\right)^{2}-n_{1}.$
@@
Refer to Corollary 5.35 in Vershynin, (2010). ∎
@@
[…]
@@
For any matrix $Z\in\mathbb{R}^{n_{1}\times n_{2}}$, let its $l$th largest singular value be $s_{l}(Z)$ for $l=1,\ldots,n\wedge p$. Then for all $1\leq l\leq n\wedge p,$
[…]
@@
he first term is bounded by $\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\|(M_{0}^{\top}M_{0}/n)-\mathbb{I}_{k}\|\lesssim 1/\sqrt{n}$ with probability at least $1-o(1)$, using Lemma E.1. The second term is handled by first observing that $\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}\mid M_{0}\sim N\left(0,\sigma_{0v}^{2}\lambda_{0u}^{\top}M_{0}^{\top}M_{0}\lambda_{0u}\right)$ and then Lemma E.1 to argue that $|\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}|\lesssim 1/\sqrt{n}$ with probability at least $1-o(1).$ The third term is seen to be $O_{P_{0}}(1/n)$ and the fourth term is seen to be $O_{P_{0}}(n^{-1}+p_{n}^{-1})$ using Proposition 3.5 along with $\|y^{(u)}\|_{2}\lesssim\sqrt{n}$, with probability at least $1-o(1)$. Therefore, one has the desired result for all $1\leq u,v\leq p$.
@@
First, consider $u,v$ such that $u\neq v$. Following from Assumption 3.2, there exists $W_{1}>0$ such that $\|\lambda_{0u}\|_{2}>c_{1}$ for all $1\leq u\leq p$. Let $c_{2}=\min_{u}\sigma_{0u}^{2}>0$ from Assumption 3.3. Following from Lemma E.4, we have for sufficiently large $n$ with probability at least $1-o(1)$: (1) $\|\mu_{u}\|_{2}^{2}>c_{1}/2$ and $\mathcal{V}_{u}^{2}>c_{2}/2,$ (2) $\left|\lambda_{0u}^{\top}\lambda_{0v}\right|\leq\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\leq k\|\Lambda_{0}\|_{\infty}^{2}$. This implies that with probability at least $1-o(1)$,
$b_{uv}^{2}\leq 1+\dfrac{4k^{2}\|\Lambda_{0}\|_{\infty}^{4}}{c_{1}c_{2}}<\infty.$
@@
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least $1-o(1)$.
@@
$\max_{1\leq j\leq p}|F_{j}|\leq\dfrac{\gamma_{0}\delta_{0}^{2}}{\gamma_{n}}+\dfrac{1}{n}\left\|P_{U_{0}}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2}+\dfrac{\gamma_{0}}{n\gamma_{n}}\left\|\mathbb{I}_{n}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2},$
as $\widehat{M}\widehat{M}^{\top}=nUU^{\top}$. Next, we have $\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2}\lesssim n$ with probability at least $1-o(1)$, following Lemma 1 in Laurent and Massart, (2000). We start with
@@
with probability at least $1-o(1)$. The second inequality is obtained using Proposition 3.5. Next, we work with
@@
$\lesssim$  $\,1$
Combining the bounds for $G_{1}$ and $G_{2}$, we have with probability at least $1-o(1),$ $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}.$
@@
with probability at least $1-o(1).$ In particular, if $k=o(n)$, we have $\|M_{0}\|\asymp\sqrt{n}$ with probability at least $1-o(1).$
@@
Appendix F Proofs of Related Lemmas for Theorems 3.7 and 3.8
@@
Suppose in the setup of Theorem 3.7, we have $T_{n}=Z_{n}+Y_{n}$ for random variables $T_{n},Z_{n},Y_{n}$ such that the pseudo-posterior of $Y_{n}$ concentrates around 00 as $n,p\to\infty$ and the pseudo-posterior of $Z_{n}$ is the $N(0,1)$ density. Then as $n,p\to\infty,$
@@
where the inequality in the second line above is obtained by observing $\Phi(x)=\Pi_{C}(Z_{n}\leq x)=\Pi_{C}(Z_{n}\leq x,Y_{n}\leq\epsilon)+\Pi_{C}(Z_{n}\leq x,Y_{n}>\epsilon)$ and then using the triangle inequality. The first term can be bounded by $\Phi(x)-\Phi(x-\epsilon)\leq L\epsilon$ for a universal constant $L>0$, as $\Phi$ is Lipschitz continuous. The second term goes to 00 in $P_{0}$-probability, and thus is less than $L\epsilon$ with probability at least $1-o(1)$ for sufficiently large $n$. Thus, for sufficiently large $n$, $S_{1}\leq 2L\epsilon$ with probability at least $1-o(1)$. Similarly, $S_{2}\leq 2L\epsilon$ with probability at least $1-o(1)$ for sufficiently large $n$. Combining, we have

### v3

@@ B.3 Proof of part (c)
$d_{uv}^{*}(\rho)=\sqrt{\dfrac{n}{n+\tau^{-2}}}d_{uv}(\rho)R_{uv}+\sqrt{n}\left(\dfrac{\rho^{2}\widetilde{\sigma}_{v}\widetilde{\sigma}_{u}\tilde{e}_{u}^{\top}\tilde{e}_{v}}{n+\tau^{-2}}\right).$
Due to FABLE-posterior concentration of $\widetilde{\sigma}_{j}^{2}$ to $\sigma_{0j}^{2}$ and convergence in probability of $\mu_{j}||_{2}^{2}$ to $\lambda_{0j}||_{2}^{2}$, we have FABLE-posterior concentration of $d_{uv}(\rho)$ around 0 as $n\to\infty$ for any finite $\rho>0$. Furthermore, with probability at least $1-o_{P_{0}}(1)$,
$\sqrt{n}\left(\dfrac{\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}\tilde{e}_{u}^{\top}\tilde{e}_{v}}{n+\tau^{-2}}\right)\lesssim\dfrac{1}{\sqrt{n}}.$
Thus, the FABLE-posterior of $d_{uv}^{*}(\rho)$ concentrates around 0 asymptotically. As $n\to\infty$, Lemma F.2 leads to
$\underset{x}{\sup}\,\left|\widetilde{\Pi}\left\{\dfrac{\sqrt{n}(\widetilde{L}_{uv}-\mu_{u}^{\top}\mu_{v})}{l_{0,uv}(\rho)}\leq x\right\}-\Phi(x)\right|\overset{P_{0}}{\to}0.$
For $u=v$, we first observe that $\sqrt{n}(\widetilde{\Psi}_{uu}-T_{uu})=\sqrt{n}(\widetilde{L}_{uu}-\|\mu_{u}\|_{2}^{2})+\sqrt{n}(\widetilde{\sigma}_{u}^{2}-\delta_{u}^{2}).$ Using the decomposition in part (b) of the proof of Theorem 3.7, we can express
$\sqrt{n}(\widetilde{\sigma}_{j}^{2}-\delta_{j}^{2})=-\dfrac{2\sqrt{n}\,\Delta_{j}}{\gamma_{n}}\sigma_{0j}^{2}+\mathcal{G}_{j},$
for all $1\leq j\leq p$, where
$\mathcal{G}_{j}=-\dfrac{2\sqrt{n}(\delta_{j}^{2}-\sigma_{0j}^{2})(\Delta_{j}/\gamma_{n})}{1+(2\Delta_{j}/\gamma_{n})}+\dfrac{4\sqrt{n}\sigma_{0j}^{2}(\Delta_{j}^{2}/\gamma_{n}^{2})}{1+(2\Delta_{j}/\gamma_{n})}.$
Letting $\mathcal{N}_{j}=-\dfrac{2\sqrt{n}\,\Delta_{j}}{\gamma_{n}}\sigma_{0j}^{2},$ we have $\mathcal{N}_{j}/\sqrt{2\sigma_{0u}^{4}}\implies N(0,1)$, using the normal approximation to the Gamma distribution. We now consider bounding each of the terms in $\mathcal{G}_{j}$. The first term is bounded as
$\left|-\dfrac{2\sqrt{n}(\delta_{j}^{2}-\sigma_{0j}^{2})(\Delta_{j}/\gamma_{n})}{1+(2\Delta_{j}/\gamma_{n})}\right|$ $\asymp|\mathcal{N}_{j}|\,|\delta_{j}^{2}-\sigma_{0j}^{2}|$
$\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}$
$=o(1)$
with probability at least $1-o_{P_{0}}(1)$, by the result on $\mathcal{N}_{j}$ and the bound on $|\delta_{j}^{2}-\sigma_{0j}^{2}|$. The second term is analogously bounded, which provides $|\mathcal{G}_{j}|\lesssim o(1)$ with probability at least $1-o_{P_{0}}(1)$. Thus, we can write
$\sqrt{n}\left(\widetilde{\Psi}_{uu}-T_{uu}\right)=\left(2\rho\,\sigma_{0u}\|\lambda_{0u}\|_{2}\right)R_{uu}+\mathcal{N}_{u}+\mathcal{G}_{u}.$
@@
and $f_{n}=n/(n+\tau^{-2})^{2}.$ Observe that $f_{n}=(1/n)+\Delta_{n}$, where $|\Delta_{n}|\asymp 1/n^{2}$. Since $y^{(u)}=M_{0}\lambda_{0u}+\epsilon^{(u)}$, we can decompose $S_{uv}$ as
@@
We first break up $S_{uv}$ into two parts $S_{uv}=D_{uv}+R_{uv},$ where
@@
We first deal with $R_{uv}$ and show that $\sqrt{n}R_{uv}=o_{P_{0}}(1)$. We use throughout that $\|M_{0}\|\asymp\sqrt{n}$ and $\|\epsilon^{(u)}\|_{2}\lesssim\sqrt{n}$ with probability at least 1o(1),1-o(1), using Lemma E.8. First consider bounding $\sqrt{n}f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}$ as follows:
@@
The first term is $\sqrt{n}f_{n}\left|\epsilon^{(u)}U_{0}U_{0}^{\top}\epsilon^{(v)}\right|\,=O_{P_{0}}(\sqrt{n}f_{n})=O_{P_{0}}(1/\sqrt{n})$, since $U_{0}^{\top}\epsilon^{(u)}\sim N(0,\sigma_{0u}^{2}\mathbb{I}_{k})$ and $k$ is finite. The second term is handled by observing that with probability at least 1o(1)1-o(1), we have
@@
    We first observe that with probability at least 1o(1)1-o(1),
@@
[…]
@@
    We next observe that with probability at least 1o(1),1-o(1),
@@
[…]
@@
$\left.\dfrac{\sqrt{n}}{n}\left(\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}\epsilon^{(u)}\right)\,\right|M_{0}\sim N(0,l_{uv}^{2}(M_{0}))\overset{d}{=}l_{uv}(M_{0})Z_{uv},$
where $Z_{uv}\sim N(0,1)$ and $Z_{uv}$ is independent of $M_{0}.$ Let
@@
Since $\|M_{0}\|\asymp\sqrt{n}$, we have $\left|l_{uv}(M_{0})+l_{0,uv}\right|=O_{P_{0}}(1)$ and $\left|l_{uv}(M_{0})+l_{0,uv}\right|\geq l_{0,uv}\geq\sqrt{2c_{1}^{2}c_{2}}>0$, from Assumptions 3.2 and 3.3. Thus, the third term can be handled by observing
@@
$=O_{P_{0}}\left(\sqrt{\dfrac{\log n}{n}}\right),$
as $\left\|\dfrac{M_{0}^{\top}M_{0}}{n}-\mathbb{I}_{k}\right\|=O_{P_{0}}\left(\sqrt{\dfrac{\log n}{n}}\right)$ from Lemma E.1. Since $Z_{uv}$ is independent of $M_{0}$, Lemma F.1 immediately implies that
$\sqrt{n}D_{uv}\overset{d}{=}N(0,\xi_{0,uv}^{2}+l_{0,uv}^{2})+o_{P_{0}}(1).$
@@
since $\|AB\|\geq\|A\|s_{\min}(B)$ for any matrices $A,B$, with $s_{\min}(B)$ denoting the smallest singular value of $B$. Both $\Lambda_{0}$ and $M_{0}$ have exactly $k$ positive singular values since the rank is $k$, so that $\min\left(s_{\min}(\Lambda_{0}),s_{\max}(M_{0})\right)>0$. Since the Schatten $\infty$-norm is identical to the spectral norm, Theorem 2 in Luo et al., (2021) implies
@@
Thus with probability at least 1o(1)1-o(1),
@@
Using Lemma E.8, we have $s_{\min}(M_{0})\asymp\|M_{0}\|\ \asymp\sqrt{n}$ with probability at least 1o(1)1-o(1), since $k=O(1)$. Under Assumption 3.2, $s_{\min}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$. Suppose $\sigma_{\text{sum}}^{2}=\sum_{j=1}^{p}\sigma_{0j}^{2}$ and $\sigma_{\max}^{2}=\max_{1\leq j\leq p}\sigma_{0j}^{2}=\mathcal{O}(1)$ by Assumption 3.3. Lemma E.2 implies $\|E\|\lesssim(\sigma_{\text{sum}}+\sqrt{n}\sigma_{\max})$ with probability at least 1o(1)1-o(1). Since $\sigma_{\text{sum}}\leq\sqrt{p_{n}}\sigma_{\max}$, one has $\|E\|\lesssim\sqrt{n}+\sqrt{p_{n}}$ with probability at least 1o(1)1-o(1), as $n\to\infty$. Thus, with probability at least 1o(1)1-o(1),
@@ E.2 Other Lemmas for Theorem 1
For the true factor matrix $M_{0}$, we have
$\mathbb{P}\left(\left\|\dfrac{M_{0}^{\top}M_{0}}{n}-\mathbb{I}_{k}\right\|\leq 3\sqrt{2}\,\sqrt{\dfrac{\log n}{n}}\right)\geq 1-\dfrac{2}{n}.$
@@
Refer to Corollary 5.35 and Lemma 5.36 in Vershynin, (2010). ∎
@@
Suppose $E_{0}\in\mathbb{R}^{n_{1}\times n_{2}}$ has independent entries such that $E_{0,ij}=g_{ij}b_{ij}$ for $g_{ij}\overset{iid}{\sim}N(0,1)$ and $\{b_{ij}:1\leq i\leq n_{1},1\leq j\leq n_{2}\}$ are fixed scalars. Let $\sigma_{1}=\underset{i}{\max}\sqrt{\sum_{j}b_{ij}^{2}}$, $\sigma_{2}=\underset{j}{\max}\sqrt{\sum_{i}b_{ij}^{2}},$ and $\sigma_{*}=\underset{i,j}{\max}|b_{ij}|.$ Then for every $\epsilon\in(0,1/2]$ there exists a $c_{\epsilon}^{{}^{\prime}}>0$ such that for all $t\geq 0$,
@@
For any matrix $Z\in\mathbb{R}^{n_{1}\times n_{2}}$, let its $l$th largest singular value be $s_{l}(Z)$ for $l=1,\ldots,n_{1}\wedge n_{2}$. Then for all $1\leq l\leq n_{1}\wedge n_{2},$
$s_{l+1}(Z)=\min_{\text{rank}(M)\leq l}\|Z-M\|.$
@@
he first term is bounded by $\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\|(M_{0}^{\top}M_{0}/n)-\mathbb{I}_{k}\|\lesssim\sqrt{(\log n)/n}$ with probability at least 1o(1)1-o(1), using Lemma E.1. The second term is handled by first observing that $\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}\mid M_{0}\sim N\left(0,\sigma_{0v}^{2}\lambda_{0u}^{\top}M_{0}^{\top}M_{0}\lambda_{0u}\right)$ and then Lemma E.1 to argue that $|\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}|/n\lesssim 1/\sqrt{n}$ with probability at least 1o(1).1-o(1). The third term is seen to be $O_{P_{0}}(1/n)$ since $U_{0}^{\top}\epsilon^{(v)}\mid U_{0}\sim N_{k}(0,\sigma_{0v}^{2}\mathbb{I}_{k})$ and the fourth term is seen to be $O_{P_{0}}(n^{-1}+p_{n}^{-1})$ using Proposition 3.6 along with $\|y^{(u)}\|_{2}\lesssim\sqrt{n}$, with probability at least 1o(1)1-o(1). Therefore, one has the desired result for all $1\leq u,v\leq p$.
@@
First, consider $u,v$ such that $u\neq v$. Following from Assumptions 3.2 and 3.3, we have $\|\lambda_{0u}\|_{2}\geq c_{1}>0$ and $\sigma_{0u}^{2}\geq c_{2}>0$ for some constants $c_{1},c_{2}$, for all $1\leq u\leq p$. Following from Lemma E.4, we have for sufficiently large $n$ with probability at least 1o(1)1-o(1): (1) $\|\mu_{u}\|_{2}^{2}>c_{1}^{2}/2$ and $\mathcal{V}_{u}^{2}>c_{2}/2,$ (2) $\left|\lambda_{0u}^{\top}\lambda_{0v}\right|\leq\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\leq k\|\Lambda_{0}\|_{\infty}^{2}$. This implies that with probability at least 1o(1)1-o(1),
$b_{uv}^{2}\leq 1+\dfrac{4k^{2}\|\Lambda_{0}\|_{\infty}^{4}}{c_{1}^{2}c_{2}}<\infty.$
@@
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least 1o(1)1-o(1).
@@
[…]
@@
with probability at least 1o(1)1-o(1). The second inequality is obtained using Proposition 3.6. Next, we work with
@@
$\lesssim$  $\,1.$
Combining the bounds for $G_{1}$ and $G_{2}$, we have with probability at least 1o(1),1-o(1), $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}.$ This completes the proof. ∎
@@
with probability at least 1o(1).1-o(1). In particular, if $k=o(n)$, we have $\|M_{0}\|\asymp\sqrt{n}$ with probability at least 1o(1).1-o(1).
@@
Appendix F Proofs of Related Lemmas for Theorems 3.8 and 3.9
@@
Suppose in the setup of Theorem 3.8, we have $T_{n}=Z_{n}+Y_{n}$ for random variables $T_{n},Z_{n},Y_{n}$ such that the posterior of $Y_{n}$ concentrates around 0 as $n,p\to\infty$ and the posterior of $Z_{n}$ is the $N(0,1)$ density. Then as $n,p\to\infty,$
@@
where the inequality in the second line above is obtained by observing $\Phi(x)=\Pi_{C}(Z_{n}\leq x)=\Pi_{C}(Z_{n}\leq x,Y_{n}\leq\epsilon)+\Pi_{C}(Z_{n}\leq x,Y_{n}>\epsilon)$ and then using the triangle inequality. The first term can be bounded by $\Phi(x)-\Phi(x-\epsilon)\leq L\epsilon$ for a universal constant $L>0$, as $\Phi$ is Lipschitz continuous. The second term goes to 0 in $P_{0}$-probability, and thus is less than $L\epsilon$ with probability at least 1o(1)1-o(1) for sufficiently large $n$. Thus, for sufficiently large $n$, $S_{1}\leq 2L\epsilon$ with probability at least 1o(1)1-o(1). Similarly, $S_{2}\leq 2L\epsilon$ with probability at least 1o(1)1-o(1) for sufficiently large $n$. Combining, we have